A boat builder uses labor and capital to produce boats as described by the following production function: Q (L, K) = 120K'L? - KL? In the short-run, the firm's capital is fixed at 10 units. a. Obtain equation for the firm's total product curve. b. Obtain the firm's average and marginal product functions


1
Expert's answer
2021-02-05T13:32:54-0500

q=120KL-KL

K=10

Q=120OL+10L

APL=Q/L

(1200L+1OL)/L

APL=1210

marginal product of labour=average product of labour

marginal product of labour

MPL=dQ/dL

mpl=1200+10